# PositionIC: Unified Position and Identity Consistency for Image Customization

### v3

@@ Introduction
Quantitative Evaluations
Subject-driven Analyses
Specifically, we discover that different object sizes can lead to inaccurate scores. As shown in Figure 6, to avoid the sensitivity, we crop the subject from original image as source images for evaluation.
Controllable Spacial Generation
@@
User Study
We invite evaluators for an extensive user study. For subject-driven generation, we randomly selected 500 images from the results on DreamBench for manual evaluation. There are six users scored the results from five dimensions, with each ranging from 0 to 5, and the average score was taken. The result presented in Figure 8 (a) shows that our method reaches the highest capability to preserve image features and details while maintaining competitive text adherence.
@@
Ablation Study
In this section we show the ablation study of our key modules, including horizon mask and data quality. The results are shown in Figure 9, Table 4 and Figure 10.
Impact of data filter.
@@
w/o horizon mask 0.784 0.269 0.686
w/ horizon mask 0.846 0.269 0.823
Table 4: Ablation study of horizon mask. Our model perform better subject fidelity on Dreambench after restrict the attention horizon of reference images.
Figure 10: Qualitative results of ablation study. Horizon mask and our filter pipeline are capable of effective position control and consistent customization.
Horizon mask.
As shown in Table 4, CLIP-I and DINO scores significantly drop when training without horizon mask. We believe that by adopting horizon mask, model can focus the transfer and generation of image features on a smaller region rather than globally, which accelerates the convergence and improve the consistency.
Conclusion
Appendix A Bidirectional Multi-dimensional Perception Data Synthesis
Detailed Instruction of GPT-4o
@@ Appendix A Bidirectional Multi-dimensional Perception Data Synthesis
Due to the substantial differences in textual descriptions between subjects, it is not feasible to predetermine the feature categories for evaluation. Therefore, we allow the LLMs to select at least three features and assign individual scores to each. The final score is calculated as the average of all feature scores. For descriptions with significant discrepancies, the LLMs is permitted to assign a score of zero to the samples.
@@
Details of PIC-98K Dataset
We propose PIC-400K utilizing our Bidirectional Multi-dimensional Perception Data Synthesis, a automatic and effective high-consistency data synthesis pipeline. Samples of the filtered data PIC-98K is shown in the Figure 13. BMPDS can synthesize high-fidelity multi-subject images while maintaining high resolution. Against previous works, the position of subjects is controllable and it is randomly placed to train the position control capability of PositionIC.
@@
Appendix B PositionIC-Bench

### v4

@@ 4.3.1 Subject-driven Analyses
4.3.2 Controllable Spacial Generation
@@
4.3.3 User Study
We invite evaluators for an extensive user study. For subject-driven generation, we randomly selected 500 images from the results on DreamBench for manual evaluation. Six users scored the results from five dimensions, each ranging from 0 to 5, and the average score was taken. The result presented in Figure 8 (a) shows that our method reaches the highest capability to preserve image features and details while maintaining competitive text adherence.
@@ 4.3.3 User Study
4.4 Ablation Study
In this section we show the ablation study of our key modules, including data quality and Volumetric Weight Regulation(VWR). More ablation results are presented in the appendix.
4.4.1 Impact of data filter.
@@ 4.4 Ablation Study
Figure 9: Impact of VWR. In the presence of overlaps, VWR is capable of correctly handling the spatial relationships between objects. We set the attention value in overlap region as 1 to disable VWR.
4.4.2 Abation results of Volumetric Weight Regulation(VWR)
As shown in Fig. 9, in Case 1 overlap setting, the doll is instructed to be placed behind the sports car. However, the model fails to correctly understand their spatial relationship without VWR. In Case 2, VWR is able to prevent the generation of missing objects caused by concept confusion, even when multiple objects overlap.
5 Conclusion
@@
Supplementary Material
Figure 10: Category distribution of PIC-98K.
In Section 6, we give a detailed description of our data synthesis and filtering pipeline: Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS). We further present a selection of example images. Section 7 introduces the PositionIC-Bench we use for position evaluation.
Section  8,9 and 10 provide a detailed discussion of specific experimental setups and additional ablation studies. Furthermore, more example demonstrations can be found in Figure 12.
6 Bidirectional Multi-dimensional Perception Data Synthesis
6.1 Detailed Instruction of GPT-4o
@@
Due to the substantial differences in textual descriptions between subjects, it is not feasible to predetermine the feature categories for evaluation. Therefore, we allow the LLMs to select at least three features and assign individual scores to each. The final score is calculated as the average of all feature scores. For descriptions with significant discrepancies, the LLMs are permitted to assign a score of zero to the samples.
@@
6.2 Details of PIC-98K Dataset
We propose PIC-400K utilizing our Bidirectional Multi-dimensional Perception Data Synthesis, an automatic and effective high-consistency data synthesis pipeline. Samples of the filtered data PIC-98K are shown in the Figure 16. BMPDS can synthesize high-fidelity multi-subject images while maintaining high resolution. Against previous works, the position of subjects is controllable and it is randomly placed to train the position control capability of PositionIC.
@@ 6.2 Details of PIC-98K Dataset
Figure 12: More results about PositionIC.
7 PositionIC-Bench
@@
8 Training Details and Result
8.1 Setting of Training
The value of the semantic density $\sigma_{i}$ can be manually set or determined by Eq. 8. In  Eq. 8, we set the semantic density to a geometric progression with a common ratio of $1+10\lambda$, which represents that $\sigma_{i}$ increases incrementally from far to near to reflect the difference between foreground and background objects.
$\sigma_{i}=\frac{10\lambda*(1+10\lambda)^{i}}{(1+10\lambda)^{n}-1},$ (8)
where $n$ is the number of reference images and $i=\{0,1,2,...,n-1\}$ represents the object index ordered from far to near. $\lambda$ is a hyperparameter used to control the variation in semantic density. A smaller value of $\lambda$ results in less difference in semantic density between objects. During training, we set it to 5.
8.2 More Result
As shown in Figure 12, we downloaded several foreground images from the internet as additional visual demonstrations, covering four categories: Characters, Items, Animals, and Space. In the Characters and Items sections, PositionIC exhibits high fidelity and is capable of generating images with coherent spatial relationships. In the Animals section, we demonstrate the plausibility of object interactions under different prompts. In the Space section, PositionIC is able to achieve good results even for difficult examples (e.g., generating a planetary ring surrounding the Earth).
9 Revised Evaluation Metric
In this section, we elaborate on the calculation methods for our evaluation metrics. As shown in Figure 13, directly using the entire image to compute the CLIP or DINO metrics is unreasonable, as both CLIP and DINO compute the similarity of global image features. To avoid the sensitivity, we crop the subject from original image as source images for evaluation.
Figure 14: Qualitative results of ablation study. Visibility-Aware Attention(VAA) and our filter pipeline are capable of effective position control and consistent customization.
10 More Ablation Study
In this section, we conduct more detailed ablation studies, including ablations on the data filter and Visibility-Aware Attention(VAA).
10.1 Impact of Visibility-Aware Attention
As shown in Figure 4, CLIP-I and DINO scores significantly drop when training without VAA. We infer that incorporating a attention mask allows the model to focus more effectively on features in a smaller region rather than globally, which accelerates the convergence and improves the consistency.
10.2 Qualitative Results
As shown in Figure 14, we visualize the effect of VAA and data filtering. It can be seen that without VAA, the model loses the ability to control position. When using unfiltered data, although the data volume is larger, the generated quality is poorer.
[…]
w/o VAA 0.784 0.269 0.686
w/ VAA 0.846 0.269 0.823
Table 4: Ablation study of VAA. Our model performs better subject fidelity on DreamBench after restricting the attention area of reference images.